Mrs. Garson received 3 new students. There are a total of 210 different ways they could sit. How many seats are available for them to sit in?
Answer by sudhanshu_kmr(1152)

let the total no. of seat = x
then no. of ways by which 3 people can sit is xP3 = 210 ( x permutation 3)
=> x!/(x-3)! = 210
=>x*(x-1)*(x-2) = 210
=> x* (x-1) *(x-2) = 2 * 3 * 5 *7 by factorizing 210
=>x*(x-1)*(x-2) = 7*6*5


so, value of x is 7.

no. of seats = 7